OpenStudy (anonymous):

Solve the following system of equations for x. - 3x + 7y = 2 - 2x - 3y = - 4

OpenStudy (lgbasallote):

if you multiply the FIRST equation by 3 and the SECOND equation by 7 you get... -9x + 21y = 6 -14x - 21y = -28 now add them up -23x = -22

OpenStudy (anonymous):

2x - 3y = 7 -3x + y = 7 2x - 3y = 7 -9x + 3y = 21 -7x = 28 x = 28/-7 = -4 -3(-4)+ y = 7 12 + y = 7 y = -5 Your solution is (-4, -5)